OpenStudy (unklerhaukus):

Use partial fractions or the convolution theorem and the table of Laplace transforms, to find functions of \(t\) which have the following Laplace transforms. (b) \[\frac{p+2}{2p^2+p-1}\]

OpenStudy (unklerhaukus):

(b) partial fractions \[\begin{align*} F(p)&=\frac{p+2}{2p^2+p-1}\\ &=\frac{p+2}{(p+1)(2p-1)}&=\frac{A}{p+1}+\frac{B}{2p-1}\\ \\&&p+2=A(2p-1)+B(p+1)\\ \\&\text{for } p=-1&1=-3A\\ &&A=-\tfrac{1}3\\ \\&\text{for } p=0&2=-A+B\\ &&B=2-\tfrac 13=\tfrac53\\ \\ F(p)&=\frac{-\tfrac13}{p+1}+\frac{\tfrac53}{2p-1}\\ &=\tfrac13\left(\frac{5}{2p-1}-\frac{1}{p+1}\right)\\ \\ f(t)&=\tfrac13\mathcal L^{-1}\left\{\frac{5}{2p-1}-\frac{1}{p+1}\right\}\\ &=\tfrac13\left(\tfrac52\mathcal L^{-1}\left\{\frac{1}{p-\tfrac12}\right\}-\mathcal L^{-1}\left\{\frac1{p+1}\right\}\right)\\ &=\tfrac13\left(\tfrac52e^{t/2}-e^{-t}\right)\\ \\ &=\frac56e^{t/2}-\frac13e^{-t}\\ \end{align*} \]

OpenStudy (unklerhaukus):

what do i do with that/??

OpenStudy (unklerhaukus):

im having trouble with deconvolution

OpenStudy (anonymous):